First Call for Papers - ICDT97

6th International Conference on

Database Theory

Delphi, Greece, January 8-10, 1997

The Sixth International Conference on Database Theory (ICDT ’97) will take place in Delphi, Greece, on January 8-10, 1997. ICDT is a biannual international conference on theoretical aspects of databases and a forum for communicating research advances on the principles of database systems. Previous ICDT conferences were held in Rome (1986), Bruges (1988), Paris (1990), Berlin (1992), and Prague (1995).

Topics of interest: ICDT aims to attract papers of high quality, describing original ideas and new results on theoretical aspects of all forms of databases. Major themes to be covered include, but are not limited to:

active database systems  complex objects  complexity and optimization

concurrency control and recovery data models and design theory data structures and algorithms for databases database programming languages datalog analysis deductive databases dependencies and constrains distributed and heterogeneous databases fundamentals of security and privacy

incomplete information  multimedia information systems  nonmonotonic reasoning
object-oriented databases  query evaluation and optimization  query languages

transaction management updates

Submissions: Authors are invited to submit 8 copies of a full conference paper (which might be an extended abstract where proofs are only sketched) of at most 5000 words (10 typed pages in no less than 11-point font) before June 1, 1996, to one of the program co-chairs:

Program co-chair:  Foto Afrati  Phokion Kolaitis

Computer Science Division Computer and Information Sciences National Technical University of Athens University of California Heroon Polytechniou 9,157 73 Zographou Santa Cruz, CA 95064 Athens, Greece USA Tel: +30-1-772-2498 Tel: +1-408-459-4768 Fax: +30-1-772-2459 Fax: +1-408-459-4829

                        afrati@softlab.ece.ntua.gr            kolaitis@cse.ucsc.edu

The cover page of each paper should contain the address, e-mail, fax and phone numbers of the contacting author as well as a short abstract. The results presented in the paper must be unpublished and not submitted for publication elsewhere, including the proceedings of other conferences or workshops. One author of each accepted paper will be expected to present the paper at the conference. As in the past, it is planned that the proceedings of ICDT’97 will be published by Springer-Verlag in the Lecture Notes in Computer Science series, and will be available at the conference.

Program Committee:

F.Afrati (Greece)  M.Cadoli (Italy)  S.Chaudhuri (USA)  S.Cluet (France)
T.Eiter (Austria)  C.Faloutsos (USA)  S.Greco (Italy)  M.Gyssens (Belgium)
P.Kanellakis (USA)  O.H.G.Katona (Hungary)  P.Kolaitis (USA)  G.Kuper (Germany)
T.Milo (Israel)  M.Ozsoyoglu (USA)  J.Pokorny (Czech Republic) V.Sazonov (Russia)
H-J.Schek (Switzerland)  E.Soisalon-Soininen (Finland) J.Su (USA)  V.Tannen (USA)
D.Van Gucht (USA)  V.Vianu (USA)  G.Vossen (Germany)  P.Wolper (Belgium)

J.Zlatuska (Czech Republic)

Organization co-chair: Foto Afrati, Theodora Varvarigou

Important dates:  Deadline for submission:  May 31, 1996

Notification of acceptance or rejection: September 3, 1996 Camera-ready papers due: October 4, 1996 Conference: January 8-10, 1997

web page: http://